OpenStudy (anonymous):

how would i solve 3^(x+4)=7^(x-1)

OpenStudy (anonymous):

logarithms

hero (hero):

\[3^{x+4}=7^{x-1}\] Expand both sides the rule \(a^{b + c} = a^ba^c\) Express the equation as a proportion in the following manner: \(\dfrac{3^x}{7^x} = \dfrac{1}{7(3^4)}\) Apply \(\dfrac{a^c}{b^c} = \left(\dfrac{a}{b}\right)^c\) to the left side: \(\left(\dfrac{3}{7}\right)^x = \dfrac{1}{7(81)}\) \(\left(\dfrac{3}{7}\right)^x = \dfrac{1}{567}\) Now take the logs of both sides: \(x \log(3/7) = \log (1/567)\) \(x = \dfrac{\log(1/567)}{\log(3/7)}\)

hero (hero):

Note that \(x\) can be simplified further to get \(x = \dfrac{\log(7) + \log(3^4)}{\log(7) - \log(3)}\)